You may need to use the appropriate appendix table or technology to answer this question.

Consider the following hypothesis test.

H0: μ1 − μ2 = 0

Ha: μ1 − μ2 ≠ 0

The following results are for two independent samples taken from the two populations.

Sample 1 Sample 2

n1 = 80

n2 = 70

x1 = 104

x2 = 106

σ1 = 8.2

σ2 = 7.4

(a)

What is the value of the test statistic? (Round your answer to two decimal places.)

(b)

What is the p-value? (Round your answer to four decimal places.)
Answer #1

SOLUTION:

From given data,

The following results are for two independent samples taken from the two populations.

Sample 1 Sample 2
n_{1} = 80 n_{2} = 70
x_{1} = 104 x_{2} = 106
\sigma _{1} = 8.2 \sigma _{2} = 7.4

The statistical software output for this problem is:

Two sample Z summary hypothesis test:

\mu _{1} : Mean of population 1 (Std. dev. = 8.2)
μ2 : Mean of population 2 (Std. dev. = 7.4)
\mu _{1} - μ2 : Difference between two means
Null hypothesis : H0 : \mu _{1} - μ2 = 0
Alternative hypothesis : Ha : \mu _{1} - μ2\neq 0

(a) What is the value of the test statistic? (Round your answer to two decimal places.)

Test statistic

z = (x_{1} - x_{2} ) / sqrt (\sigma_{1}^{2} / n_{1} + \sigma_{2}^{2} / n_{2})

z = (104 - 106 ) / sqrt (8.22 / 80 + 7.42 / 70)

z = -2 / 1.273886

z = -1.56

(b) What is the p-value? (Round your answer to four decimal places.)

p-value

p-value = P (z > -1.56 )

= 1 - P (z < -1.56 )

= 1 - 0.05938

= 0.94062

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.